OpenStudy (anonymous):

A cube with volume of 8 cubic cm is inscribed in a circle so that each vertex of the cube touches the sphere. What is the length of the diameter, in centimeters, of the sphere? Please explain process...

OpenStudy (anonymous):

sphere... the cube is inscribed in a sphere

OpenStudy (unklerhaukus):

the diameter of the sphere will be the length of the diagonal through the cube, you will want to find the dimensions of the cube first

OpenStudy (unklerhaukus):

to find the side length of the cube use \[V_{cube}=s^3=8cm^3\]

OpenStudy (anonymous):

so the edge length of the cube would be 2

OpenStudy (anonymous):

how would you find the diagonal of the cube? I am a plane thinker, a plane slicing the cube in three vertices would create the diagonal as the hypotenuse.

OpenStudy (anonymous):

so that would make the diagonal = 3.46?

OpenStudy (anonymous):

\[2^{2}+2^{2}=8\rightarrow8+4=12\rightarrow \sqrt{12}=3.46\]

OpenStudy (anonymous):

look right?

OpenStudy (unklerhaukus):